OpenStudy (anonymous):

the width of a rectangle is 0.6 of the length. if the width is 9 ft what is the length? the width of a rectangle is 0.6 of the length. if the width is 9 ft what is the length? @Mathematics

OpenStudy (anonymous):

W=0.6*L 9=0.6*L 9/0.6=L

OpenStudy (anonymous):

 Isolate for L that is the Length